I'm having trouble. I solved for y not x. It was:
2x+y=5
1x-2y=4
I ended up doing this:
2x+y=5
-2x -2x
________
y=5-2x
etc. what did i do wrong?

Answer by edjones(8007)

2x+y=5
1x-2y=4
.
4x+2y=10 multiply each side by 2
x-2y=4
--------add
5x=14
x=14/5
.
14/5 -2y = 4
14-10y=20 multiply each side by 5
-10y=6
y=-6/10=-3/5
.
Ed